Solve the following equations and check your results.8x + 4 = 3 (x – 1) + 7 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

8x+4=3(x–1)+7
⇒8x+4=3x–3+7
⇒8x+4=3x+4
⇒8x-3x=4-4
⇒5x=0
⇒x=0
Putting the value of x in RHS and LHS we get,
(8×0) + 4 = 3 (0–1) + 7
⇒0+4=0-3+7
⇒4=4
⇒LHS=RHS

Thanks